# Instructions for ACL 2023 Proceedings

### v1

@@ 3.1 VLM Dataset Curation
ICC 500k 14.6 34.9 47.0 30.6 60.9 72.9
Table 2: Representation learning results over different filtered dataset sizes. We perform text-to-image retrieval evaluation over MS-COCO and Flickr for different filtering schemes of LAION-400M with varying dataset sizes. We compare our performance (ICC) to various filtering baselines: Rand. indicates selecting random samples from LAION-400M, CLIP indicates filtering by top CLIPScore, CA indicates Complexity and Action filtering, and TMS indicates filtering with T-MARS. Best results are in bold.
We also perform a representation learning experiment by training a dual text and image encoder model on LAION-400M filtered with different methods. Table 2 reports text-to-image retrieval over standard held-out retrieval benchmarks, namely MS-COCO Lin et al. (2014) and Flickr Plummer et al. (2015). The model is initialized from pretrained vision and text encoders (vit-base, BERT-BaseDosovitskiy et al. (2010); Devlin et al. (2018), as suggested by Zhai et al. (2022). We use a batch size of 128, learning rate of 2e-5 with a cosine scheduling function. All other filtering methods in the table are identical to the ones in the captioning model training. As illustrated in the table, ICC yields superior performance for this task, showing that our method selects samples which provide better signal for downstream retrieval applications.
We note that although prior work has found filtering methods such as CLIPScore to be beneficial, we find that it fails to significantly improve (or even degrades) results in the case of selecting a small core of samples. This accords with previous work showing that applying filtering to LAION-400M with CLIP degrades the performance Maini et al. (2023) in some of the benchmarks, likely due to high-scoring images containing literal text that overlaps with the caption.
@@ 3.2 Concreteness Correlation
Word Conc. Sentence Conc.
@@
ICC 0.75 0.75 0.55 0.69 0.67 0.54
Table 3: Concreteness evaluation on single-word and sentence-level texts. Correlation is measured using Pearson $\rho$, Spearman $\rho_{s}$, and Kendall $\tau$ coefficients.
Table 3 shows the correlations of different concreteness estimation methods to ground-truth concreteness scores on both single-word and sentence-level (caption) benchmarks. We compare to zero-shot probing of CLIP through Stroop probing (SP) as proposed by Alper et al. (2023). We also compare to aveCLIP Wu and Smith (2023), the only (to the best of our knowledge) learned metric quantifying concreteness at the sentence level, which generates multiple images from a caption and measures the average CLIP-similarity between the text and generated images. Due to its high computational cost, we only evaluate it on a statistically-significant portion of the single-word benchmark, which contains nearly 15K samples.
Correlation to Word Concreteness. We first validate our metric by measuring it on a dataset introduced by Hessel et al. (2018). This dataset is composed of 39,954 English unigrams and bigrams coupled with human-labelled concreteness scores on a scale from 1 (abstract) to 5 (concrete), averaged over annotators. To compare with prior work, we only use unigram nouns, totaling 14,562 items. As illustrated in Table 3, ICC significantly outperforms prior works over all correlation metrics.
Correlation to Caption Concreteness. We manually annotated concreteness scores for 200 captions from LAION-400M Schuhmann et al. (2022), selected to cover a wider variety of levels of concreteness compared to random samples from the same dataset, by following a rule-based method. We provide further details of the sampling method used in the appendix. We emphasize that these captions are different from the ones used to find the optimal combinations of the VBA and SBA, thus avoiding data leakage. As seen in Table 3, ICC outperforms existing methods in this setting as well, by an even larger margin than that of the single-word setting.
Sentence Concreteness
Pipe Model Part Sim. $\rho$ $\rho_{s}$ $\tau$
SBA TinyLLaMa-1.1B ED 0.49 0.43 0.33
SBA OpenLLaMa-3B ED 0.48 0.44 0.35
SBA LLaMa-2-7B ED 0.50 0.42 0.31
SBA TinyLLaMa-1.1B BSc 0.45 0.41 0.32
SBA OpenLLaMa-3B BSc 0.45 0.40 0.31
SBA LLaMa-2-7B BSc 0.47 0.42 0.32
VBA BLIP-Base ED 0.47 0.41 0.32
VBA BLIP-Large ED 0.39 0.29 0.23
VBA BLIP-2 ED 0.41 0.34 0.26
VBA BLIP-Base BSc 0.42 0.45 0.35
VBA BLIP-Large BSc 0.54 0.48 0.37
VBA BLIP-2 BSc 0.55 0.47 0.37
Table 4: Ablations over VBA and SBA Design Choices. We ablate the effect of the LLM used in the SBA pipeline and the captioning model used in the VBA pipeline, as well as the sentence-similarity measure, on the correlation to the ground-truth concreteness annotations. Note that here we measure correlation to each model of the piplines (VBA and SBA) used in isolation. BSc and ED refer to BERTScore and edit distance respectively. We report the Pearson $\rho$, Spearman $\rho_{s}$, and Kendall $\tau$ correlation coefficients. Best results are in bold, and second-best are underlined
@@ 4 Ablations
ICC Model Parts Ablation. In Table 4, we ablate the effect of design choices in the ICC pipeline, namely different LLM sizes  Zhang et al. (2024); Geng and Liu (2023) in the SBA pipeline and different captioning model architectures in the VBA, as well as the use of edit distance or BERTScore as the similarity measure in each pipeline, in order to justify the use of LLaMa-7B and edit distance in the SBA and BLIP2 and BERTScore in VBA. Note that we ablate the correlation to ground-truth annotations of each of the pipelines in isolation. As can be seen, increasing the size of the LLM in the SBA results in a mild increase in correlation to the ground-truth annotations. Similarly, a more capable captioning model in the VBA provide similar improvement. Moreover, we can also see that the edit distance similarity measure produces better correlation in the SBA pipeline, while BERTScore fits better in the settings of the VBA. This observation matches the intuition that the SBA reconstructs captions with more ease, benefiting from a more stringent metric for detecting abstract versus concrete reconstructions as opposed to the VBA, which should rely on a more semantic measure to properly detect abstract sentences.
$\rho$ $\rho_{s}$ $\tau$
Before Distillation 0.60 0.53 0.42
After Distillation 0.69 0.67 0.54
Table 5: Distillation Effect on Caption-Concreteness Correlation. We show correlations to ground-truth annotated caption concreteness scores before and after distillation. The “After Distillation" row corresponds to our final ICC score.
Distillation Concreteness Effect. Although the distillation procedure is necessary to make inference feasible with respect to runtime, we provide further motivation by measuring the effect of distillation on the correlation to ground-truth annotations of concreteness scores in Table 5. As can be seen, the distillation improves the correlations, providing further motivation beyond computational efficiency and simplifying the inference of our ICC model. We hypothesize that this improvement is due to smoothing of noisy reconstruction of the VBA and SBA by the distillation process.
Distillation Speedup. We ablate the speed-up provided by the distillation phase by running the SBA, VBA and the distilled ICC on the same hardware settings (an Nvidia A6000 GPU), the same batch size of 1 and the same caption samples. We find that the SBA and VBA process 0.45 and 0.2 samples per second respectively, and the distilled score processes 45 samples per second. Note that the time it would take to generate scores for our 8M subset of LAION-400M dataset is approximately 11,000 GPU hours for the VBA and 5,000 GPU hours for the SBA compared to just 50 GPU hours using the distilled ICC. Additionally, for a batch size of 1, the distilled model takes less than 700 MB of GPU memory compared to 13GB and 14GB for the VBA and SBA respectively.
Use of Both SBA and VBA Scores. We also conduct an ablation of the importance of using both the SBA and VBA scores for downstream captioning model training in Table 6. In the figure, we show captioning metrics (CIDEr and SPICE) of a model trained on a distilled version of each one of the scores in isolation, compared to the combined ICC metric which outperforms both.
@@
ICC 30.9 9.7 28.2 6.0
Table 6: Score Ablations We ablate the importance of using scores obtained from both the SBA and VBA pipelines over 200k samples filtered dataset using the different scores.
@@ 5 Related Work
Evaluating Text Concreteness. Word concreteness is a topic of interest in cognitive science Schwanenflugel (2013), and a number of works have studied automatic prediction of word concreteness using machine learning Hill et al. (2014); Hill and Korhonen (2014); Hessel et al. (2018); Rabinovich et al. (2018); Charbonnier and Wartena (2019); Alper et al. (2023). However, little attention has been paid to measuring concreteness at the sentence or string level. Most similar to us is Wu and Smith (2023), who generate multiple images for each caption and average the CLIP similarity scores over all the images to produce a sentence-level concreteness score. Other text evaluation metrics compare to reference texts Gehrmann et al. (2023) or a reference image Hessel et al. (2021), while we are interested in the inherent quality of text in isolation (namely, its visual concreteness).
@@ 6 Conclusion
We present a new metric for measuring the visual concreteness of image captions without an image reference. By leveraging strong foundation models, we quantify visual-semantic information loss and find that this highly correlates with human concreteness judgments. Our results demonstrate that ICC is effective at selecting a core of high-quality image-caption samples from web-scale multimodal datasets for training models in the resource-constrained setting. We foresee the use of ICC in additional tasks requiring the curation of web-scale multimodal data, where high-quality, visually-concrete text is needed.
@@
While our method manages to detect visually concrete captions well, it lacks sensitivity to grammatical structure, which might cause it to label oddly phrased captions as concrete. For instance, consider the caption: “a computer near a tree with a boy next to a table with a keyboard”. This caption is highly concrete and gets a high ICC score of 1.0. However, removing all object relations from the caption produces the following: “computer tree boy table keyboard” which results in a relatively minor decrease of the ICC score to 0.89. Such low-quality captions might have a negative impact on tasks such as image captioning where the model must learn to output grammatically correct English sentences which should ideally describe relevant fine-grained relations between entities. We hypothesize that this behavior stems from the dataset used to train the distillation model (CC3M) which is not likely to include such oddly phrased caption, and so these non-grammatical structures are not learned. We hypothesize that training over a dataset with higher caption diversity will likely alleviate this issue.
In addition, due to limited computational resources, our experiments were conducted on a relatively small scale of 8 million sample dataset. We expect that increasing the scale and the filtered dataset proportionally will result in a performance improvement in the downstream model performance. However, we leave verifying this as well as testing the effect of ICC filtering on other downstream tasks such as VQA and caption ranking to future work.
@@ A.1 Standardizing By Caption Length
We aim to have reconstruction scores that are only dependent on the concreteness of captions and not on the length of the captions. In Figure 6, we show the distribution of the reconstruction similarities before and after standardization per caption length. We can see in Figure 5(a) that there is a strong dependency on caption length, which we would like to avoid.
@@ A.2 ICC Distillation
We distill the knowledge obtained by the two pipelines described in the paper in a two-stage manner. Firstly, we distill the VBA and SBA scores into two distinct DistilRoBERTa Liumm et al. (2019) models. We then collect a small subset of 244 captions, sampled to have approximately uniform joint distribution of scores, and annotate the concreteness scores of these captions. This is showcased in Figure 4. We regress over these samples to get the optimal weights as discussed in the main paper.
(a) Random sampling
(b) Our test dataset
Figure 7: Distribution of annotated concreteness scores of 200 random samples (6(a)) and our test dataset (6(b)) sampled using rule-based methods. All samples are from LAION-400M. 0 is highly in-concrete while 3 means highly concrete caption.
A.3 Caption Concreteness Benchmark
Next we describe the data collection and annotation details for the test set used in the main paper to compute the correlations to human concreteness intuition. Our aim is to have a small, yet diverse set of samples that represent the wide diversity of possible captions. Since Laion-400M is very noisy and only a small portion of it includes highly concrete captions, we sample 150 items that satisfy the following rules:
  • The caption must include at least 10 characters
  • The caption must not contain more the 80% of capitalized words.
  • The caption must include at least 2 stop words, filtered using NLTK parser Loper and Bird (2002).
  • The ratio of stop words to all the words in a captions must not exceed 20%.
The remaining 50 samples in our dataset are selected randomly to include more “raw” captions as well. For all captions in our benchmark, we also apply NSFW filtering and make sure the caption do not include offensive or personal content. We show the distribution of concreteness scores in Figure 7 compared to a subset of random samples we annotated, with the same size. As illustrated in the figure, we were able to achieve a better coverage of highly concrete captions, more than doubling the amount of highly concrete captions, using our simple rule-based sampling. We also show all the captions in our dataset in Figure 11, sorted according to the annotated concreteness scores. The annotations are in the range 003333, when 00 denotes highly abstract caption and $4$ is a highly concrete caption.
@@ A.4 Zero-Shot CLIP Concreteness Score
We adapt the Stroop Probing method Alper et al. (2023) that is originally designed to assess the concreteness of words, to captions. We follow the same procedure used when measuring concreteness of words, but replace the empty slot in the prompts with a caption rather than a single word, and use only the prompts that fit the context of caption in the black spot (i.e., we don’t use the captions “Alice giving the [*] to Bob” and “Bob giving the [*] to Alice” as they aren’t appropriate when using a caption to replace the empty slot [*]).
A.5 aveCLIP Word Concreteness
Since aveCLIP requires generating many images per word, we found that running aveCLIP over the entire word concreteness dataset is not feasible due to runtime constraints. Therefore, we sample 150 words from the dataset, and verified that it is statistically significant by measuring the p-values of the different statistical coefficients, which were all approximately 0.
A.6 Training Hyperparameters and Additional Information
SBA. We train the linear layer of the SBA with a batch-size of 128, learning rate of 2e-3 with cosine scheduler and a warm-up ratio of 0.03, and train for a two epoch over a single Nvidia-A6000 GPU. All other hyperparameters are set to the default of HuggingFace Trainer.
VBA Text-To-Image Hyperparameters For the image generation in VBA, we use guidance scale of 9 and 20 inference steps.
CA
Tiger cubs playing in the rain at the Zoo of the Ozarks in Washington, D.C. on Saturday, Oct. 18
Coffee at the bar. I love this place! It’s a great way to get away from the hustle and bustle
how to install window blinds in your home
Cambodia, the largest of all the African savannahs, is one of the most arid regions in the world.
Rugby World Cup 2019: The men’s singles final takes place at the Ritz-Carlton in London, England, on Saturday,
TMS
Catching a lion in the wild is one of the most beautiful things you can do in the wild.
Coffee at the bar. Photo credit: Flickr userfairy.com.au. (via Flickr)
Door and window repair ideas for your home interior design ideas on pinterest
Aerial view of a wild boar in a field in Namibia, South Africa, Africa. (Photo courtesy of the Namibian Wildlife)
Astonishingly, there was no shortage of competition between the two-school teams at the London 2012 Olympic Games.
CLIP
Polar bear cubs pose for a photo with a polar bear in the background. Credit: NASA/JPL-Caltech/UCLA
Pizza Hut Creamy Pizza Sandwich with Bacon, Cheese, and Tomato Sauce
Poster for the new ""Polar Bear"" on the front of the house. It features a picture of a polar bear with the words
Aerial view of the world’s largest crocodile in the Serengeti National Park.
New York Mets Fanatics Authentic 8"" x 10"" Skateboard Deck
ICC
Panda eating bamboo
A picture of a pizza box full of pizzas.
Outdoor Glass Window
Zebra at the zoo
A view of the tennis court from the front.
Figure 8: More qualitative examples of captions generated by captioning models trained on datasets filtered with different filtering methods, over images from MS-COCO validation split. T-MARS is marked by TMS.
A.7 Model Checkpoints Used

### v4

@@ 3.1 VLM Dataset Curation
PAC 500k 8.8 23.5 33.9 17.8 40.4 53
ICC 500k 14.6 34.9 47 30.6 60.9 72.9
Table 2: Representation learning results over different filtered dataset sizes. We perform text-to-image retrieval evaluation over MS-COCO and Flickr30K for different filtering schemes of LAION-400M with varying dataset sizes. We compare our performance (ICC) to various filtering baselines: Rand. indicates selecting random samples from LAION-400M, CLIP indicates filtering by top CLIPScore, CA indicates Complexity and Action filtering, TMS indicates filtering with T-MARS and PAC indicates filtering with PACScore. Best results are in bold.
Image-Text Representation Learning. We also perform a representation learning experiment by training a dual text and image encoder model on LAION-400M filtered with different methods. Table 2 reports text-to-image retrieval over standard held-out retrieval benchmarks, namely MS-COCO Lin et al. (2014) and Flickr30K Plummer et al. (2015). The model is initialized from pretrained vision and text encoders (ViT-base, BERT-BaseDosovitskiy et al. (2010); Devlin et al. (2018), as suggested by Zhai et al. (2022). We use a batch size of 128, learning rate of 2e-5 with a cosine scheduling function. All other filtering methods in the table are identical to the ones in the captioning setting. As illustrated in the table, ICC yields superior performance for this task, showing that our method selects samples which provide better signals for downstream retrieval applications in this setting.
We note that although prior work has found filtering methods such as CLIPScore Hessel et al. (2021) to be beneficial Gadre et al. (2023), we find that it fails to significantly improve (or even degrades) results in the case of selecting a small core of samples. This accords with previous work showing that applying filtering to LAION-400M with CLIP degrades the performance Maini et al. (2023) in some of the benchmarks, likely due to high-scoring images containing literal text that overlaps with the caption.
@@ 3.2 Concreteness Correlation
Word Conc. Caption Conc.
@@
GPT-3.5 0.55 0.56 0.44 0.44 0.48 0.4
GPT-4o 0.78 0.79 0.64 0.57 0.57 0.49
ICC 0.75 0.75 0.55 0.73 0.75 0.6
Table 3: Concreteness evaluation on single-word and caption-level texts. Correlation (in absolute value) is measured using Pearson $\rho$, Spearman $\rho_{s}$, and Kendall $\tau$ coefficients. Best result are in bold, second best are underlined.
Table 3 shows the correlations of different concreteness estimation methods to ground-truth concreteness scores on both single-word and caption-level benchmarks. We compare ICC to three baselines. The first baseline is zero-shot probing of CLIP through Stroop probing (SP) as proposed by Alper et al. (2023). The second baseline is aveCLIP Wu and Smith (2023), a learned metric quantifying concreteness at the sentence level, which generates multiple images from a caption and measures the average CLIP-similarity between the text and generated images. Due to its high computational cost, we evaluate it on a random subset of each benchmark (as described in the appendix). Finally, we compare to GPT-3.5-Turbo and GPT-4o Achiam et al. (2023), used in the zero-shot setting by prompting them to provide concreteness scores. The prompts used are detailed in the appendix.
Correlation to Word Concreteness. We first validate our metric by measuring it on the dataset introduced by Hessel et al. (2018). This consists of 39,954 English unigrams and bigrams coupled with human-labelled concreteness scores on a scale from 1 (abstract) to 5 (concrete), averaged over annotators. To compare with prior work, we only use unigram nouns, totaling 14,562 items. As seen in Table 3, ICC outperforms prior dedicated methods for measuring word concreteness, while performing competitively with the proprietary and much larger GPT-4o.
Correlation to Caption Concreteness. We manually annotate concreteness scores for 500 captions from LAION-400M Schuhmann et al. (2022), selected to cover a wide variety of levels of concreteness. As seen in Table 3, ICC outperforms existing methods in this setting by a large margin, demonstrating its advantage in selecting the most concrete image captions.
$\rho$ $\rho_{s}$ $\tau$
Before Distillation 0.65 0.6 0.46
After Distillation 0.72 0.75 0.6
Table 4: Distillation Effect on Caption Concreteness Correlation. We show correlations to ground-truth annotated caption concreteness scores before and after distillation. The “After Distillation" row corresponds to our final ICC score.
@@ 4 Ablations
Distillation Concreteness Effect. Although the distillation procedure is necessary to make inference feasible with respect to runtime, we provide further motivation by measuring the effect of distillation on the correlation to ground-truth annotations of concreteness scores in Table 4. As can be seen, the distillation improves correlations values, providing further motivation beyond computational efficiency and simplifying the inference of our ICC model. We hypothesize that this improvement is due to smoothing of noisy reconstruction of the VBA and SBA by the distillation process.
Distillation Speed-up. We ablate the speed-up provided by the distillation phase by running the SBA, VBA and the distilled ICC on the same hardware settings (an Nvidia A6000 GPU), the same batch size of 1 and the same caption samples. We find that the SBA and VBA process 0.45 and 0.2 samples per second respectively, and the distilled score processes 45 samples per second. Note that the time it would take to generate scores for our 8M subset of LAION-400M dataset is approximately 11,000 GPU hours for the VBA and 5,000 GPU hours for the SBA compared to just 50 GPU hours using the distilled ICC. Additionally, for a batch size of 1, the distilled model takes less than 700 MB of GPU memory compared to 13GB and 14GB for the VBA and SBA respectively.
Use of Both SBA and VBA Scores. We also ablate the use of both SBA and VBA scores for downstream captioning model training in Table 5. In the figure, we show captioning metrics (CIDEr and SPICE) of a model trained on a distilled version of each of the scores in isolation, compared to the combined ICC metric which outperforms both.
@@
ICC 30.9 9.7 28.2 6
Table 5: Score Ablations We ablate the importance of using scores obtained from both the SBA and VBA pipelines over 200k samples dataset that was filtered using the different scores.
ICC Model Component Ablations. In Table 6, we ablate the effect of various design choices in the ICC pipeline by evaluating their effects on caption concreteness prediction (using the benchmark described in Section 3.2). In particular, we test different LLM sizes  Zhang et al. (2024); Geng and Liu (2023) in the SBA pipeline, different captioning model architectures in the VBA pipeline, and the similarity measure used in each pipeline (edit distance vs. BERTScore). To identify the effect of each component, we evaluate SBA and VBA predictions in isolation (without combining or distilling them). As is seen in the table, our chosen LLM and captioning model perform comparably to the alternative models tested, showcasing the robustness of the VBA and SBA across model sizes. Moreover, while the simple edit distance similarity measure performs acceptably for the SBA pipeline, the BERTScore similarity measure produces significantly better correlations in the VBA pipeline, matching the intuition that the VBA is inherently lossy with respect to the precise form of texts and must rely on a more semantic measure to properly detect abstract sentences.
Caption Concreteness
Pipe Model Part Sim. $\rho$ $\rho_{s}$ $\tau$
SBA TinyLLaMa-1.1B ED 0.59 0.58 0.45
SBA OpenLLaMa-3B ED 0.57 0.56 0.43
SBA LLaMa-2-7B ED 0.53 0.51 0.48
SBA TinyLLaMa-1.1B BSc 0.57 0.56 0.43
SBA OpenLLaMa-3B BSc 0.56 0.55 0.42
SBA LLaMa-2-7B BSc 0.57 0.56 0.43
VBA BLIP-Base ED 0.43 0.4 0.31
VBA BLIP-Large ED 0.43 0.36 0.27
VBA BLIP-2 ED 0.44 0.41 0.31
VBA BLIP-Base BSc 0.6 0.6 0.46
VBA BLIP-Large BSc 0.58 0.56 0.43
VBA BLIP-2 BSc 0.6 0.58 0.45
Table 6: Ablations over VBA and SBA Design Choices. We ablate the effect of the LLM used in the SBA pipeline and the captioning model used in the VBA pipeline, as well as the text similarity measure, on the correlation to the ground-truth concreteness annotations. Note that here we measure correlation to each model of the piplines (VBA and SBA) used in isolation. BSc and ED refer to BERTScore and edit distance respectively. We report the Pearson $\rho$, Spearman $\rho_{s}$, and Kendall $\tau$ correlation coefficients. Our default settings are indicated with a prepended *.
@@ 5 Related Work
Evaluating Text Concreteness. Word concreteness is a topic of interest in cognitive science Paivio et al. (1968); Richardson (1975); Schwanenflugel (2013); Khanna and Cortese (2021), and a number of works have studied automatic prediction of word concreteness using machine learning Hill et al. (2014); Hill and Korhonen (2014); Hessel et al. (2018); Rabinovich et al. (2018); Charbonnier and Wartena (2019); Alper et al. (2023). However, little attention has been paid to measuring concreteness at the caption or string level. Shi et al. (2019) define concreteness of constituents by matching them to images for learning syntactic representations without explicit supervision; as was later shown, the signal of noun concreteness plays a key role in the model’s syntactic predictions Kojima et al. (2020). Most similar to us is Wu and Smith (2023), who generate multiple images for each caption and average the CLIP similarity scores over all the images to produce a caption-level concreteness score. Other text evaluation metrics compare to reference texts Gehrmann et al. (2023) or a reference image Hessel et al. (2021), while we are interested in the inherent quality of text in isolation (namely, its visual concreteness).
@@ 6 Conclusion
We present a new metric for measuring the visual concreteness of image captions without an image reference. By leveraging strong foundation models, we quantify visual-semantic information loss in an unsupervised manner and find that this highly correlates with human concreteness judgments. Our results demonstrate that ICC is effective at selecting a core of high-quality image-caption samples from web-scale multimodal datasets for training models in the resource-constrained setting. We foresee the use of ICC in additional tasks requiring the curation of web-scale multimodal data, where high-quality, visually-concrete text is needed.
@@
While our method manages to detect visually concrete captions well, it lacks sensitivity to grammatical structure, which might cause it to label oddly phrased captions as concrete. For instance, consider the caption: “a computer near a tree with a boy next to a table with a keyboard”. This caption is highly concrete and gets a high ICC score of 1.0. However, removing all object relations from the caption produces the following: “computer tree boy table keyboard” which results in a relatively minor decrease of the ICC score to 0.89. Such low-quality captions might have a negative impact on tasks such as image captioning where the model must learn to output grammatically correct English sentences which should ideally describe relevant fine-grained relations between entities. We hypothesize that this behavior stems from the dataset used to train the distillation model (CC3M) which is not likely to include such oddly phrased captions, and so these non-grammatical structures are not learned. We hypothesize that training over a dataset with higher caption diversity will likely alleviate this issue.
In addition, due to limited computational resources, our experiments were conducted on a relatively small scale of 8 million sample initial training dataset based on LAION-400M. We expect that increasing the scale and the filtered dataset proportionally will result in a performance improvement in the downstream model performance. However, we leave verifying this as well as testing the effect of ICC filtering on other downstream tasks such as VQA and caption ranking to future work.
@@
Acknowledgements
This work was partially supported by the KLA Foundation and Google. We thank Yonatan Bitton and Keren Ganon for their helpful feedback.
@@ A.1 Standardizing By Caption Length
We aim to have reconstruction scores that are only dependent on the concreteness of captions and not on the length of the captions for both the SBA and VBA. In Figure 5, we show the distribution of the edit-distance based reconstruction similarities of the SBA before and after standardization per caption length. We can see in Figure 5(a) that there is a strong dependency on caption length, which we would like to avoid.
@@ A.2 ICC Distillation
We distill the knowledge obtained by the two pipelines described in the paper in a two-stage manner. Firstly, we distill the VBA and SBA scores into two distinct DistilRoBERTa Liumm et al. (2019) models. We then collect a small subset of 244 captions, sampled to have approximately uniform joint distribution of scores, and annotate the concreteness scores of these captions. This is showcased in Figure 6. We regress over these samples to get the optimal weights.
Figure 7: Distribution of annotated concreteness scores in our manually labeled test set of 500 captions. All samples are from LAION-400M. Annotations range from highly abstract (0) to highly concrete (3).
A.3 Caption Concreteness Benchmark Distribution
Our aim is to have a small, yet diverse set of samples that represent the wide diversity of possible captions. Since Laion-400M is very noisy and only a small portion of it includes highly concrete captions, we curate our captions to achieve a balanced distribution of concreteness scores, as illustrated in Figure 7. As seen there, the concreteness of the benchmark’s captions is evenly distributed between abstract and concrete concepts.
@@ A.4 Zero-Shot CLIP Concreteness Score
We adapt the Stroop Probing method Alper et al. (2023) for estimating text concreteness. While Alper et al. (2023) test this on single words, we adapt this method to captions by replacing the empty slot in prompts with a caption rather than a single word. We use their prompts, omitting those which do not match the context of an entire caption being inserted in the masked slot (i.e., omitting the prompts “Alice giving the [*] to Bob” and “Bob giving the [*] to Alice”).
A.5 GPT Prompts
The following prompts were used to extract concreteness scores for image captions [^4] To get the concreteness scores of words, we used the same prompts with “word” instead of “caption” in the appropriate places. from GPT-3.5 and GPT-4o:
System: ‘‘You are an expert visual reasoner, capable of understanding the visual concretess of image captions. A visually concrete caption is a caption that is highly visual, and can be vividly imagined.’’
User: ‘‘Provide a numerical score on a scale of 1-5, when 1 is non-visual and 10 is highly visual caption for the following caption : <caption>. Only provide the numerical score and nothing else.’’
Note that we experimented with three different ranges of [1-N] of concreteness scores in our prompts: N=3, N=5 and N=10. We found that N=5 yielded the best results.
A.6 aveCLIP Word Concreteness
Since aveCLIP requires generating many images per word or caption, we found that running aveCLIP over the entire word concreteness dataset is not feasible due to runtime constraints. Therefore, we evaluate its performance on a random subset of 150 words/captions.
A.7 Training Hyperparameters and Additional Information
SBA. We train the linear layer of the SBA with a batch-size of 128, learning rate of 2e-3 with cosine scheduler and a warm-up ratio of 0.03, and train for a two epoch over a single Nvidia-A6000 GPU. All other hyperparameters are set to the defaults of the HuggingFace Trainer API.
VBA Text-To-Image. For the image generation of the diffusion model in the VBA, we use guidance scale of 9 and 20 inference steps.
A.8 Model Checkpoints Used